Solve for x and y:
x^2-4x+y^2+6y=12

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Solve for x and y:
x^2-4x+y^2+6y=12
-----------
That's the equation of a circle.
You can pick a value for x and find y, or vice versa.
There are an infinite # of points.

Answer by ikleyn(52824)

.
If to consider your post from the Math point of view then . . .

    this part "Solve for x and y" is not relevant to this part "x^2-4x+y^2+6y=12".


    And vice versa: This part "x^2-4x+y^2+6y=12" is not relevant to this ""Solve for x and y".

As a consequence, nobody understands you.


Watch your Math language !!